40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6055     Entry time: Wed Nov 30 22:09:20 2011
Author: Zach 
Type: Update 
Category: RF System 
Subject: some final EOM stabilization efforts 

First, things that were done:

  • I was troubled by the odd-looking noise in the EOM temperature signal, so I investigated the circuit with a probe and found that there was quite a bit of line pickup, which I traced to the wires going to and from the RTD (if I placed a dummy resistor directly on the board, it went down markedly).
    • I put a 3-Hz RC LPF between the AD620 and the driver input buffer, which reduced the line noise significantly
    • The error signal looks much cleaner and there are no longer strong peaks in the error spectrum at ~1+ Hz and harmonics
  • I had tried earlier to increase the gain of the servo at the driver input stage. It seemed to stay stable. Since I knew the error signal  with the loop closed was at the level of the ADC noise, I decided to push my luck with increasing the servo gain and juice up the AD620 gain from 100 to 990.
    • The servo stayed stable and the error signal level is now manageable.

Things that I noticed:

  • With the latest increase in gain, I measured that the error signal was suppressed with the loop closed (the suppression is below ~0.1 Hz, and the reason that the high-frequency level is different is because it has been amplified above the ADC noise by the time of the second trace).

 EOM_temp_stab_and_unstab_new_11_30_11.pdf

  • Despite the above, the Stochmon signals remained unchanged no matter what I did. I noticed that the Stochmon signals, too, were fluctuating basically at bit-level. I terminated the 11-MHz signal and compared it to the normal level---it is not exactly the same, but only a factor of 2-3 lower, which is not great. Of course, the RMS detector is logarithmic, but I think we still want the dark noise to be at least an order of magnitude lower here.
    • I tried to amplify the signal with an SR560, but since the DC level is supposed to be ~1-2 V, I could only get about 2x gain---not enough.

 11MHz_wandwo_ctrl_and_adc_noise_11_30_11.pdf

Conclusion

I think there are two things that could be happening here, given the above information:

  1. We are limited by the noise of the temperature sensing circuit, which would explain why the in-loop error signal is suppressed while the RAM levels appear not to be. This should be easy enough to test (though there's not enough time right now) with an out-of-loop sensor.
  2. The RAM is not dominated by temperature noise here. With the loop open, one would expect to see coherence between the RAM signal and the temperature sensor, if indeed one was the cause of the other. Instead, we see that---while the 11- and 55-MHz signals ARE pretty coherent with each other---there is no appreciable coherence between the temperature and the Stochmon signal.
ELOG V3.1.3-